OpenStudy (anonymous):

Please help!!!!!! Is there a property that, for real numbers a and b \[\Large \frac{a}{b}\]

OpenStudy (anonymous):

\[\Large a=\frac{1}{b}~~~~=~~b=\frac{1}{a}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Just wondering is there a name to that property?

OpenStudy (anonymous):

I don't understand the question,

OpenStudy (anonymous):

You see how you can wirte a=1/b and then that is the same thing as b=1/a? Well is there a name give to that? @Jack174

OpenStudy (dumbcow):

i would call it the reciprocal property....by taking reciprocal of both sides the equality still holds

OpenStudy (dumbcow):

btw i like pi ...lol

OpenStudy (anonymous):

Ok, thanks, and no I'm concerining the pie that has to do with gastronomy, not mathematics. Sure, pi is transcedental, but that doesn't mean we have to discriminate. Lol

OpenStudy (anonymous):

And thanks,

OpenStudy (dumbcow):

haha yw